OpenStudy (anonymous):

Xavier's dining room table was unstable because one leg was shorter than the others. He placed a three-fourths inch thick piece of wood under the short leg, and it became even with the other legs, which are nineteen and one-eighth inches long. How many inches long was the short leg before Xavier added the wood? eighteen and three-eighths inches eighteen and five-eighths inches eighteen and three-fourths inches seventeen and five-eighths inches

OpenStudy (anonymous):

@amistre64 help me pls if u can

OpenStudy (amistre64):

Xavier's dining room table was unstable because one leg was shorter than the others. He placed a three-fourths inch thick piece of wood under the short leg, and it became even with the other legs, [the other legs] are nineteen and one-eighth inches long. How many inches long was the short leg before Xavier added the wood? assuming a flat floor, then: x + 3/4 = 19 + 1/8

OpenStudy (amistre64):

x = 19 + 1/8 - 3/4

OpenStudy (anonymous):

Ok let me solve

OpenStudy (anonymous):

@amistre64 Its 18.375 which is 18 and three eight inches

OpenStudy (amistre64):

18 and 3/8 is good

OpenStudy (anonymous):

Thank you for helping!

OpenStudy (amistre64):

youre welcome